OpenStudy (anonymous):

*~*PLEASE HELP!*~* What is the equation of the line in standard form that passes through the point (3, 4) and is parallel to the line y = –x + 6?

OpenStudy (anonymous):

y = –x + 6 Slope = -1 since the lines are parallel the slopes are = nw using pont slope form y-y1=m(x-x1) (3, 4) y-4 = -1(x-3) y - 4 = -x + 3 y = -x + 7